OpenStudy (icalibear):

What conjecture can you make about the sum of 10 positive even numbers?

OpenStudy (klimenkov):

This sum is even.

OpenStudy (icalibear):

2 = 2 = 1(3) 2 + 4 = 6 = 2(3) 2 + 4 + 6 =12 = 3(4) 2 + 4 + 6 + 8 = 20 = 4(5) 2 + 4 + 6 + 8 + 10 = 30 = 5(6)

OpenStudy (icalibear):

(a) The sum is 9(10) (b) The sum is 10(10) (c) The sum is 11(12) (d) The sum is 10(11)

OpenStudy (klimenkov):

What will the tenth term in this sequence be?

OpenStudy (klimenkov):

The tenth addend.

OpenStudy (icalibear):

10(11)?

OpenStudy (klimenkov):

10*10

OpenStudy (klimenkov):

OOPS.

OpenStudy (klimenkov):

10*11

OpenStudy (icalibear):

lol ok. thnks